OpenStudy (anonymous):

the length of rectangle exceeds its width by 5 inches and the area is 84 square inches. what are the length and width of the rectangle

OpenStudy (shaik0124):

A=l*b l=b+3 A=54 solve now

OpenStudy (anonymous):

3,6

Mehek (mehek14):

@candice15 no if the length is 6 and the width is 3, then the area would be 18 not 54